OpenStudy (anonymous):

for what vaules of x is the rational experession below undefined: x+5/3x^2-3 1, -3, -1, -5, 3, or 5

TheSmartOne (thesmartone):

Heya there!! @sar12389 @~ \(\Huge\cal\color{blue}{Welcome~to~Open~Sudy!!}\) Since you are new here feel free to look at this tutorial for new Open Studiers!!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

TheSmartOne (thesmartone):

\[\frac{ x+5 }{ 3x^2 -3}\] Is this your question? @sar12389 @~

TheSmartOne (thesmartone):

Oh oops again I tagged the wrong user... @bbeatsbianca @~ Is that your question and again. Welcome to OS!!

OpenStudy (anonymous):

yes that is

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

ok :D

TheSmartOne (thesmartone):

So to get an equation undefined the denominator has to equal to 0 because \(\Huge\frac{0}{x}=0\) and \(\Huge\frac{x}{0}=undefined\)

OpenStudy (anonymous):

oh okay that makes more sense! I thought so but im just confusing myself.

TheSmartOne (thesmartone):

So to find what makes the equation undefined we just set the denominator equal to 0!!

OpenStudy (anonymous):

so would the answers be 1 and -1

TheSmartOne (thesmartone):

So lets see

TheSmartOne (thesmartone):

\( 3x^2-3=0\) \( 3x^2=3\) \( x^2=1\) \( x= \pm1\) So yes!! Good job!

OpenStudy (anonymous):

awesome thank you very much!

OpenStudy (anonymous):

\[\frac{ x-1 }{ (x-3)(x+1)} \]

OpenStudy (anonymous):

what value of x is the rational expression equal to zero? @TheSmartOne

TheSmartOne (thesmartone):

In this case the numerator has to equal to 0.

TheSmartOne (thesmartone):

@bbeatsbianca @~

OpenStudy (anonymous):

so i can just plug in the numbers they give me and make the numerator 0 and that will be my answer?

TheSmartOne (thesmartone):

You make the numerator equal to 0 so x-1=0 and solve for x :D

OpenStudy (anonymous):

x=1

TheSmartOne (thesmartone):

Exactly

OpenStudy (anonymous):

thank you! @TheSmartOne

TheSmartOne (thesmartone):

No problem!!